OpenStudy (anonymous):

Perry can fill a sorter with oranges from the conveyer belt in 10 minutes. While Perry fills the sorter, Gillian takes oranges out of the sorter and puts them in shipping crates. With Gillian taking oranges out of the sorter, it takes 25 minutes for Perry to fill the sorter. Which of the following can be used to determine the amount of time it takes for Gillian to empty the sorter if Perry does not add oranges?

OpenStudy (anonymous):

1 over 10 minus 1 over x equals 1 over 25 1 over 10 minus 1 over 25 equals 1 over x 1 over 25 minus 1 over x equals 1 over 10 1 over 10 minus 1 over x equals x over 25

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok another one for the books i think they want you to solve via \[\frac{1}{20}+\frac{1}{x}=\frac{1}{25}\] that is my best guess can you provide me with the name of the person who wrote this?

OpenStudy (anonymous):

sorry, typo i meant \[\frac{1}{10}+\frac{1}{x}=\frac{1}{25}\]

OpenStudy (anonymous):

Thanks you:0

OpenStudy (anonymous):

Try Xavier high school loll

OpenStudy (anonymous):

Wait, I think it's B?

OpenStudy (anonymous):

i think it is A

OpenStudy (anonymous):

but like i said i wouldn't bet on it

OpenStudy (anonymous):

1/10 - 1/x =1/25 ?

OpenStudy (anonymous):

Nah, I believe you :)

OpenStudy (anonymous):

actually i think you might be correct

OpenStudy (anonymous):

Thanks so much for all your help!

OpenStudy (anonymous):

Don't doubt yourself! I think you're right, now that I look at the question.

OpenStudy (anonymous):

if you solve \[\frac{1}{10}+\frac{1}{x}=\frac{1}{25}\] you get a negative answer of \(-\frac{50}{3}\) whereas if you solve with a minus sign you get \(\frac{50}{3}\) so it depends on whether you want a negative of positive answer my guess is B now too